OpenStudy (anonymous):

Three #s, a,b, and c, form an arithmetic sequence. Increasing a by 1 or c by 2 results in a geometric sequence. Find b, with an explanation please!

OpenStudy (anonymous):

a = a b = a+d c = a + 2d Geometric: a+1 b c+2 (a+1)(c+2) = b^2 substitute b = a+d and c = a+2d in to this equation

OpenStudy (anonymous):

thank you! explained it better than my teacher

OpenStudy (anonymous):

actually it says "or" so you should get (a+1)(c) = b^2 and (a)(c+2) = b^2 so you can solve for both "a" and "d"